1List the properties of homothetic and linearly homogeneous production function

1
Expert's answer
2021-10-18T11:28:51-0400
  • Homothetic production functions have the property that "f(x)=f(y)" implies "f(\lambda x)=f(\lambda y)."
  • Homogeneous production functions have the property that "f(\lambda x)=\lambda^kf(x)" for some value k.
  • Homogeneity of degree one implies constant returns to scale.
  • Homogeneous implies homothetic, but not conversely. For instance, "f(x_1,x_2)=x_1x_2+1" is homothetic but not homogeneous.
  • A function is homogeneous if it is homogeneous of degree "\alpha" for some "\alpha" "\epsilon" R.
  • A function, f is linearly homogeneous if it is homogeneous of degree 1. Along any ray from the origin, a homogeneous function defines a power function. If f is linearly homogeneous, then the function defined along any ray from the origin is a linear function.